OpenStudy (anonymous):

A rocket travels 40 km vertically upward at a 90° angle from point A to point B. It then moves a distance of 40 km along the line segment BC

OpenStudy (anonymous):

What is the height of the rocket, at point C, from the ground?

OpenStudy (anonymous):

40 km 60 km 70 km 80 km

OpenStudy (anonymous):

sin 30 = Perpendicular/40

OpenStudy (anonymous):

sin 30 = 1/2 Perpendicular = 20km Height of Rocket = 40+20 = 60 Km

OpenStudy (anonymous):

can you help with another ?

OpenStudy (anonymous):

post it

OpenStudy (anonymous):

The moon forms a right triangle with the Earth and the sun during one of its phases, as shown below http://www.photoshop.com/users/doctorwho15/assets/e59cc76ea4e24f4fad6c80a51e8ff06e A scientist measures the angle x and the distance y between the Earth and the sun. Using complete sentences, explain how the scientist can use only these two measurements to calculate the distance between the moon and the sun